Hey guys, im stuck on this word problem and was wondering if someone could show me step by step on how to solve this equation. Im confused when it comes to word problems, Thanks in advance
A marine biologist wants to set up n aquarium containing 3% salt water. He has two tank on hand that contain 6% and 2% salt water. How much water from each tank must he use to fill a 32-gallon aquarium with a 3% saltwater mixture?

Answer by nerdybill(7384)

A marine biologist wants to set up n aquarium containing 3% salt water. He has two tank on hand that contain 6% and 2% salt water. How much water from each tank must he use to fill a 32-gallon aquarium with a 3% saltwater mixture?
.
Let x = amount (gallons) of 6% salt water
then
32-x = amount (gallons) of 2% salt water
.
our equation:
.06x + .02(32-x) = .03(32)
.06x + 0.64 -.02x = 0.96
.04x + 0.64 = 0.96
.04x = 0.32
x = 8 gallons of 6% salt water
.
amount of 2% salt water:
32-x = 32-8 = 24 gallons